### v3

@@ 4.2. Tropical invariant and $F$-invariant
(i) We write $[{\bf g}]=\{{\bf g}_{u}^{t}=(g_{1;u}^{t},\ldots,g_{m;u}^{t})^{T}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and $[{\bf g}^{\prime}]=\{{\bf g}_{u^{\prime}}^{t}=(g_{1;u^{\prime}}^{t},\ldots,g_{m;u^{\prime}}^{t})^{T}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$. Assume by contradiction that the two tropical points $[{\bf g}]$ and $[{\bf g}^{\prime}]$ are not sign-coherent. Then there exist some vertex $s\in\mathbb{T}_{n}$ and some $k\in[1,n]$ such that $g_{k;u}^{s}\cdot g_{k;u^{\prime}}^{s}<0$. Without loss of generality, we can assume $g_{k;u}^{s}<0$ and $g_{k;u^{\prime}}^{s}>0$.
@@
Using the chart ${\bf x}_{s}$ to calculate the $F$-invariant $(u\mid\mid u^{\prime})_{F}$, we have
@@
which contradicts $(u\mid\mid u^{\prime})_{F}=0$. Hence, the two tropical points $[{\bf g}]$ and $[{\bf g}^{\prime}]$ must be sign-coherent.
(ii) For any vertex $t$, consider the Laurent expansions of $u$ and $u^{\prime}$ with respect to vertex $t$.
@@
where $H_{uu^{\prime}}^{t}:=F_{u}^{t}F_{u^{\prime}}^{t}$ is a polynomial in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$.
By (i) and Lemma 4.26, we know that $[{\bf g}\uplus{\bf g}^{\prime}]:=\{{\bf g}_{u}^{t}+{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ forms a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. Then by the definition of good elements, we know that the product $uu^{\prime}$ is a $[{\bf g}\uplus{\bf g}^{\prime}]$-good element in $\mathcal{U}$. ∎
4.3. Log-canonicality and $F$-compatibility
In this subsection, we discuss the link between the compatible Poisson structure of (upper) cluster algebra and the $F$-compatibility.
As before, we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\Lambda)$, and let $\mathcal{U}$ be the upper cluster algebra associated to the cluster pattern $\mathcal{S}_{X}$. Denote by $S=diag(s_{1},\ldots,s_{n})$ the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, namely, $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$, where ${\bf 0}$ is the $n\times(m-n)$ zero matrix.
[…]
$\{x_{i;t_{0}},x_{j;t_{0}}\}:=\lambda_{ij;t_{0}}\cdot x_{i;t_{0}}x_{j;t_{0}}.$
@@ 4.3. Log-canonicality and $F$-compatibility
[…]
@@
Clearly, any two cluster variables in the same cluster are log-canonical. In Theorem 4.29 (ii) we will give the converse result.
@@
where $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$. If $u$ and $x_{k;w}$ are log-canonical for some $k\in[1,n]$, then $f_{k}=0$, where $f_{k}$ is the highest degree of $y_{k}$ in $F$.
@@
[…]
@@
Assume by contradiction that the highest degree $f_{k}$ of $y_{k}$ in $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}$ is non-zero. Then the left side of (4.1) is non-zero, which implies $c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k}\neq 0$. Now we compare the constant term on both sides of (4.1). The constant term on the left side is 00 while the constant term on the right side is $(c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot 1=c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k}\neq 0$. This is a contradiction. Hence, $f_{k}=0$. ∎
@@
    Let $k$ be an integer in $[1,n]$ and let $({\bf x}_{t},\widetilde{B}_{t})=\mu_{k}({\bf x}_{w},\widetilde{B}_{w})$. Suppose that $u$ and $x_{k;w}$ are not log-canonical but $u$ and $x_{i;w}$ are log-canonical for any $i\in[1,n]\backslash\{k\}$, then $u$ is a cluster monomial in ${\bf x}_{t}$.

### v4

@@ 4.2. Tropical invariant and $F$-invariant
(i) We write ${\bf g}]=\{{\bf g}_{u}^{t}=(g_{1;u}^{t},\ldots,g_{m;u}^{t})^{T}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and ${\bf g}^{\prime}]=\{{\bf g}_{u^{\prime}}^{t}=(g_{1;u^{\prime}}^{t},\ldots,g_{m;u^{\prime}}^{t})^{T}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$. Assume the contrary that the two tropical points $\bf g$}] and $\bf g$$^{\prime}$] are not sign-coherent to deduce a contradiction. Then there exist some vertex $s\in\mathbb{T}_{n}$ and some $k\in[1,n]$ such that $g_{k;u}^{s}\cdot g_{k;u^{\prime}}^{s}<0$. Without loss of generality, we can assume $g_{k;u}^{s}<0$ and $g_{k;u^{\prime}}^{s}>0$.
@@
By using the cluster ${\bf x}_{s}$ to calculate the $F$-invariant $(u\mid\mid u^{\prime})_{F}$, we have
@@
which contradicts $(u\mid\mid u^{\prime})_{F}=0$. Hence, the two tropical points $\bf g$}] and $\bf g$$^{\prime}$] must be sign-coherent.
(ii) For any vertex $t$, consider the Laurent expansions of $u$ and $u^{\prime}$ with respect to vertex $t$:
@@
where $H_{uu^{\prime}}^{t}\coloneqq F_{u}^{t}F_{u^{\prime}}^{t}$ is a polynomial in $\mathbb{Z}_{\geq 0}[y_{1},\ldots,y_{n}]$ with constant term $1$.
By (i) and Lemma 4.23, we know that ${\bf g}\uplus{\bf g}^{\prime}]\coloneqq\{{\bf g}_{u}^{t}+{\bf g}_{u^{\prime}}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ forms a tropical point in $\mathcal{S}_{Y}(\mathbb{Z}^{\rm max})$. Then by the definition of good elements, we know that the product $uu^{\prime}$ is a ${\bf g}\uplus{\bf g}^{\prime}]-$good element in $\mathcal{U}$. ∎
4.3. Log-canonical cluster variables are compatible
In this subsection, we show that if two cluster variables are log-canonical in a $\mathsf{\Lambda}$-upper cluster algebra, then they are contained in the same cluster.
As before, we fix a Langlands-Poisson triple $(\mathcal{S}_{X},\mathcal{S}_{Y},\mathsf{\Lambda})$, and let $\mathcal{U}$ be the corresponding $\mathsf{\Lambda}$-upper cluster algebra. Denote by $S=diag(s_{1},\ldots,s_{n})$ the type of the compatible pair $(\widetilde{B}_{t_{0}},\Lambda_{t_{0}})$, namely, $\widetilde{B}_{t_{0}}^{T}\Lambda_{t_{0}}=(S\mid{\bf 0})$, where $\bf 0$} is the $n\times(m-n)$ zero matrix.
Recall that we can define a Poisson bracket $\{-,-\}$ on the ambient field $\mathbb{F}=\mathbb{Q}(x_{1;t_{0}},\ldots,x_{m;t_{0}})$ using the $m\times m$ skew-symmetric matrix $\Lambda_{t_{0}}=(\lambda_{ij;t_{0}})$ at the rooted vertex $t_{0}$:
$\{x_{i;t_{0}},x_{j;t_{0}}\}\coloneqq\lambda_{ij;t_{0}}\cdot x_{i;t_{0}}x_{j;t_{0}}.$
@@ 4.3. Log-canonical cluster variables are compatible
Two elements $f_{1},f_{2}$ in $\mathbb{F}$ are said to be log-canonical with respect to the Poisson bracket $\{-,-\}$, if
@@
Clearly, any two cluster variables in the same cluster are log-canonical. In Theorem 4.26 (ii), we show the converse statement is also true.
@@
where $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}$ is a polynomial in $\mathbb{Z}[y_{1},\ldots,y_{n}]$ with constant term $1$. If $u$ and $x_{k;w}$ are log-canonical for some $k\in[1,n]$, then $f_{k}=0$, where $f_{k}$ is the maximal degree of $y_{k}$ in $F$.
@@
We know that $\Lambda_{w}=(\lambda_{ij;w})$ is the coefficient matrix of $\{-,-\}$ with respect to the cluster ${\bf x}_{w}$. Since $(\widetilde{B}_{w},\Lambda_{w})$ is a compatible pair of type $S$, we have $\widetilde{B}_{w}^{T}\Lambda_{w}=(S\mid{\bf 0})$. For $i\in[1,n]$ and $j\in[1,m]$, we denote by ${\bf e}_{i}$ the $i$th column of $I_{n}$ and $\widetilde{\bf e}_{j}$ the $j$th column of $I_{m}$. We have
@@
Assume by contradiction that the maximal degree $f_{k}$ of $y_{k}$ in $F=\sum_{{\bf v}\in\mathbb{N}^{n}}c_{{\bf v}}{\bf y}^{{\bf v}}$ is non-zero. Then the left side of (4.1) is non-zero, which implies $c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k}\neq 0$. Now we compare the constant term on both sides of (4.1). The constant term on the left side is 0 while the constant term on the right side is $(c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k})\cdot 1=c-{\bf g}^{T}\Lambda_{w}\widetilde{\bf e}_{k}\neq 0$. This is a contradiction. Hence, $f_{k}=0$. ∎
@@
[…]